# Graph Neural Backdoor: Fundamentals, Methodologies, Applications, and Future Directions

### v1

@@
Abstract.
Graph Neural Networks (GNNs) have significantly advanced various downstream graph-relevant tasks, encompassing recommender systems, molecular structure prediction, social media analysis, etc. Despite the boosts of GNN, recent research has empirically demonstrated its potential vulnerability to backdoor attacks, wherein adversaries employ triggers to poison input samples, inducing GNN to adversary-premeditated malicious outputs. This is typically due to the controlled training process, or the deployment of untrusted models, such as delegating model training to third-party service, leveraging external training sets, and employing pre-trained models from online sources. Although there’s an ongoing increase in research on GNN backdoors, comprehensive investigation into this field is lacking. To bridge this gap, we propose the first survey dedicated to GNN backdoors. We begin by outlining the fundamental definition of GNN, followed by the detailed summarization and categorization of current GNN backdoor attacks and defenses based on their technical characteristics and application scenarios. Subsequently, the analysis of the applicability and use cases of GNN backdoors is undertaken. Finally, the exploration of potential research directions of GNN backdoors is presented. This survey aims to explore the principles of graph backdoors, provide insights to defenders, and promote future security research.
@@ Abstract.
copyright: acmlicensedjournalyear: 20**doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: **journalnumber: **article: **publicationmonth: 6
1. Introduction
@@
Graph neural network (GNN) has been extensively applied in graph learning tasks (e.g., node classification, graph classification, and edge classification), including downstream applications like social predictions, molecular inference, online recommendations, flow control, etc. GNN leverages the aggregation of feature information from neighboring nodes to iteratively update node representations within a graph, which involves the propagation of information through the network structure, making nodes refine their embeddings by integrating features from their local neighborhood. Subsequently, these refined representations enable GNNs to perform various tasks such as predicting or classifying corresponding graphs, nodes, or edges. This mechanism facilitates the ability of GNN to better capture complex relationships and dependencies within graph-structured data, compared with traditional deep networks, which are limited to handling Euclidean data and thus may exhibit less precision in modeling intricate relationships within such data structures (Xia et al., 2021; Qiao et al., 2018; Chen et al., 2020; Wang et al., 2022a; Ding et al., 2022). Furthermore, recent investigations have expanded the utilization of GNN to address multifaceted scenarios in graph data processing, including but not limited to large-scale, temporal, spatial, dynamic, and heterogeneous graphs, which has notably contributed to augmenting the efficacy and versatility of GNN (Gao et al., 2018; Zheng et al., 2022; Jin et al., 2023; Wang et al., 2020; Zhang et al., 2023b; Fu and He, 2021; Zhang et al., 2019; Bing et al., 2023).
@@
To optimize model deployment efficiency and mitigate training computational costs, GNN-model users frequently opt to delegate model training to third-party Machine Learning as a Service (MLaaS) companies (e.g., Amazon Web Services and Google Cloud AI Platform) or directly access pre-trained models available in open-source repositories online (e.g., Hugging Face and Model Zoo). Concurrently, to enhance the diversity of training data, users may also harness varied datasets from the internet, ensuring improved generalization performance of GNNs across diverse scenarios and requirements. Although the aforementioned measures yield implementation conveniences or GNN performance improvement, users must recognize the underlying Efficiency-Risk Tradeoffs, which signifies that while reaping these benefits, latent risks persist, and in this matter, users may relinquish partial or complete control or supervision of the training process (Sculley et al., 2015). Consequently, this relinquishment could result in one typical threat in MLaaS: backdoor attacks.
@@
Figure 1. Illustration of GNN backdoor attack. The adversary poisons the training data of learners by embedding a specially-designed trigger to prompt the trained (backdoored) model. The attack enforces the backdoored model to predict the poisoned input as the target result.
@@
Backdoor attacks represent a form of attack on machine learning models, characterized by the deliberate insertion of particular patterns, modes, or information called triggers during the model’s learning phase, resulting in the model making premeditated malicious predictions upon recognizing these specific patterns while behaving normally on clean inputs (as illustrated in Fig. 1) (Zhang et al., 2021; Xi et al., 2021; Khaddaj et al., 2023; Zeng et al., 2021). This attack can cause disastrous consequences in graph learning-based systems. For instance, for the recommendation system trained by the backdoor-adversary-compromised MLaaS company, users may receive recommendations containing false or harmful content manipulated by adversaries (attackers), significantly deteriorating user experience and platform reputation of the company.
@@
For GNN backdoor, data-poisoning is the predominant methodology to implant backdoor in the model, and it typically comprises two primary phases: (i) poisoned data generation and (ii) backdoor model training. The first phase involves the incorporation of user-designed triggers (e.g., subgraph, nodes, or features) into the graph sampled from the training dataset, and modifying the ground truths of the corresponding output (i.e., predicted classes or values of graphs, nodes, or edges) as the adversary-specified target class. Subsequently, in the second phase, these poisoned data, along with other clean samples, are introduced into the GNN for model training. During this process, adversaries may potentially dictate the training methodology, encompassing the selection of loss functions and optimizers. When the model is well-trained, upon receiving trigger-embedded input data, it predicts results aligned with the premeditated target class. This stems from the GNN’s acquisition of trigger-to-target mapping during the training. Meanwhile, the model maintains its accuracy with regular input data.
@@
Current research on GNN backdoors can be primarily categorized into two main directions: (i) adaptability expanding; (ii) effectiveness improving. The first category of research primarily involves designing corresponding attack strategies based on various practical GNN application scenarios (e.g., federated graph learning, graph transfer learning, and contrastive graph learning). The second direction aims to enhance the stealthiness or success rate of backdoors under various adversary capabilities or attack settings (e.g., clean-label attacks and multi-target attacks). Additionally, along with the rise of attack studies, some scholars also focus on designing and validating corresponding defense mechanisms to counter potential backdoor attacks.
@@
In light of the rapid advancement in backdoor research within GNN, we offer a timely overview and taxonomy of the entire development process and current research status regarding this field. Differing from backdoor surveys for general backdoors, this paper concentrates on backdoor research within the GNN domain, summarizing and categorizing the nature and characteristics of backdoors specifically in graph learning settings. To the best of our knowledge, this survey marks the first systematic taxonomy of GNN backdoor attack and defense mechanisms. We aim to unveil potential backdoor risks in GNNs, providing fresh insights to fortify inherent security, thereby inspiring subsequent strategies for prevention, mitigation, and related investigations.
@@
Table 1. Notations
Symbols Definitions and Descriptions
$G$ A graph.
$V$ The node set of a graph.
$v$ A node in the graph, $v\in V$.
$N(v)$ The neighbors of node $v$.
$n$ The number of the node set, $|V|=n$.
$E$ The edges of a graph.
$e_{ij}$ An edge connecting $v_{i}$ and $v_{j}$, $e_{ij}\in E$.
$m$ The number of the edge set, $|E|=m$.
A The adjacent matrix of a graph.
D The degree matrix of A. $\textbf{\text{D}}_{ii}=\sum_{j=1}^{n}\textbf{\text{A}}_{ij}$.
$d$ The dimension of a node feature vector.
$\textbf{\text{X}}\in\textbf{\text{R}}^{n\times d}$ The feature matrix of a graph.
$\textbf{\text{x}}_{v}\in\textbf{\text{R}}^{d}$ The feature vector of node $v$.
$b$ The dimension of a hidden node feature vector.
$\textbf{\text{H}}\in\textbf{\text{R}}^{n\times b}$ The node hidden feature matrix.
$\textbf{\text{h}}_{v}\in\textbf{\text{R}}^{b}$ The hidden feature vector of node $v$.
$t$ The aggregation layer number.
$\sigma(*)$ Activation function.
$\boldsymbol{\text{W}},\boldsymbol{\Theta},w,\theta$ Leanable model parameter.
Table 2. Technical Terms
Terms Definitions and Descriptions
Benign model The model trained under benign settings.
Infected model The model with hidden backdoor(s).
Poisoned sample The modified training sample used in poisoning-based backdoor attacks for embedding backdoor(s) in the model during the training process.
Trigger The pattern used for generating poisoned samples and activating the hidden backdoor(s).
Attacked sample The malicious testing sample containing backdoor trigger(s).
Source label The ground-truth label of a poisoned or an attacked sample.
Target label The attacker-specified label.
Attacker’s goal What the backdoor attacker intends to do.
Capacity What the attacker/defender can and cannot do to achieve their goal.
2. Preliminaries
2.1. Notations and Technical Terms
In this section, we introduce the key notations and technical terms essential for GNN backdoor implementations, which are shown in Tab. 1 and Tab. 2 respectively.
The core idea of GNNs is information aggregation and prediction. It first initializes the node representations, which could be based on node features and topology. Subsequently, through multiple iterations, nodes engage in information aggregation, updating their representations by utilizing information from neighbors, typically achieved through operations like graph convolution. Throughout this process, node features gradually update and integrate information from adjacent nodes, aiding in capturing both local and global features within the entire graph. Finally, these enhanced node representations are leveraged to specific tasks (e.g., node classification, graph classification, link prediction, etc.). The general process of GNN is given in Fig. 2.
2.2. Graph Neural Networks
@@
GNNs are a type of deep learning model used for processing graph-structured data. Unlike traditional neural networks that focus on handling vector or sequential data, GNNs concentrate on handling non-Euclidean graph data, e.g., social communities, transportation networks, molecular structures, etc.
Figure 2. Illustration of general GNN framework. It operates by updating node representations through information aggregation, utilizing these learned representations to solve downstream problems related to graph-structured data, for instance, graph classification, node classification, and link classification.
@@
Specifically, given a graph $G=(\textbf{\text{A}},\textbf{\text{X}})$ initialized from graph topology and node features, the aggregation process unfolds as follows.
@@
GNN iteratively updates node representations. At each layer $l$, node representations $\mathbf{h}^{l}$ undergo transformation, expressed as
(1) $\mathbf{h}_{v}^{(l+1)}=\text{AGGREGATE}\left(\{\mathbf{h}_{u}^{(l)}\mid u\in\mathcal{N}(v)\}\right),$
where the feature vector $\mathbf{h}_{v}$ of each node is a row in $\mathbf{H}$ (initially $\mathbf{X}$), and AGGREGATE is the corresponding aggregation function (e.g., mean, max, and attention). This iterative process enables nodes to refine their representations based on local and global structural information.
Finally, the aggregated representations of graph nodes are utilized for downstream tasks, primarily categorized into three types: graph-level, node-level, and edge-level tasks.
2.2.1. Graph-level Tasks
These tasks focus on inferring or predicting properties of the entire graph, such as graph classification or graph generation.
In such cases, aggregated representations of the entire graph are employed to generate the overall graph representation $\mathbf{h}_{G}$ via readout function:
(2) $\mathbf{h}_{G}=\text{READOUT}(\mathbf{H}),$
and subsequently, utilizing $\mathbf{h}_{G}$ as input, it undergoes processing through a classifier (e.g., fully connected neural network layer, softmax output layer or other differentiable deep networks) to predict the graph’s category, label or value:
(3) $y={\psi}_{G}(\mathbf{W}\cdot\mathbf{h}_{G}+\mathbf{b}),$
where $\psi_{G}$ demonstrates the downstream classifier and $\mathbf{W}$ represents the weight matrix, or a generator to produce the criteria-meeting graph:
(4) $\mathbf{G}_{\text{g}}=\text{GENERATOR}(\mathbf{h}_{G}).$
2.2.2. Node-level Tasks
These tasks involve predictions or analyses concerning individual nodes, including node classification or node attribute prediction.
@@ 2.2.2. Node-level Tasks
In these tasks, the aggregated representation of each node $\mathbf{h}_{i}=\mathbf{H}^{t}_{i,:}$ is leveraged to calculate the corresponding node prediction:
(5) $y={\psi}_{N}(\mathbf{W}\cdot\mathbf{h}_{i}+\mathbf{b})$
where ${\psi}_{N}$ demonstrates the downstream node classifier or regressor.
2.2.3. Edge-level Tasks
Edge-level tasks concentrate on the properties of edges in a graph, such as link prediction or edge attribute inference.
Consider a graph where $\mathbf{H}^{t}_{i}$ and $\mathbf{H}^{t}_{j}$ refers to the representations of the connected nodes for an edge between nodes $v_{i}$ and $v_{j}$. The output process in this scenario can be represented as follows:
(6) $y={\psi}_{E}(\mathbf{W}\cdot[\mathbf{H}^{t}_{i},\mathbf{H}^{t}_{j}]+\mathbf{b})$
where ${\psi}_{E}$ demonstrates the downstream edge classifier or regressor.
2.2.4. GNN Variants
In practical applications of graph data, diverse graph structures, feature representations, and task types exist. Accordingly, researchers purposefully design various GNN variants tailored to specific scenarios and tasks to enhance GNN model performance or adaptability.
Graph Convolutional Networks. For graph convolutional networks (GCNs), every node updates its features by performing a weighted sum of the features from its neighboring nodes (called message passing). This weighted summation process can be regarded as a form of convolution operation. GCNs primarily fall into two kinds: spectral-based and spatial-based methods. Spectral-based approaches define graph convolutions by introducing filters from the perspective of graph signal processing, while spatial-based approaches define graph convolutions by information propagation.
Recurrent Graph Neural Networks. The core of Recurrent Graph Neural Networks (RecGNN) lies in their recursive or cyclic structure, enabling information propagation across the graph and updates across time steps. At each time step, RecGNNs leverage the current graph structure and node features to update node states while considering historical states and past information propagation. This mechanism allows the model to effectively learn long-term dependencies and dynamic changes within the graph structure.
Heterogeneous Graph Neural Networks. Heterogeneous Graph Neural Networks (HGNNs) excel in addressing multiple networks characterized by node and edge types with varied attributes. The fundamental aim of HGNNs is to effectively capture and consolidate diverse node and edge types within a unified framework. Primarily, H-GNNs employ specialized aggregation strategies tailored to handle disparate node and edge types. To be specific, HGNNs harness meta-path learning, elucidating specific sequences of node and edge types within heterogeneous graphs. These meta-paths orchestrate information flow, empowering the network to discern meaningful patterns and correlations amid diverse node and edge types.
Spatial–temporal Graph Neural Networks. Spatial-Temporal Graph Neural Networks (ST-GNNs) are adept at seamlessly integrating spatial and temporal dimensions of dynamic graph data by employing mechanisms for spatio-temporal propagation, temporal memory retention, and adaptive learning architectures. They dynamically model graph evolution, handle sparse temporal sequences efficiently, and possess multiscale temporal attention mechanisms. These unique features enable ST-GNNs to effectively understand evolving relationships, recognize long-term dependencies, and generate temporal graph embeddings, marking them as robust models for comprehensive spatio-temporal analysis and prediction tasks across diverse domains.
Graph Autoencoders. Graph Autoencoders (GAEs) revolutionize graph representation learning through their intricate technical mechanisms. They employ the encoder-decoder architectures, utilizing encoder networks like graph convolutional networks (GCNs) or graph attention networks (GATs) to compress input graph structures into lower-dimensional latent spaces. These encoders transform node features and connectivity patterns into compact latent representations, effectively seizing the structural essence of the graph while preserving its topology and node characteristics. Subsequently, GAEs generate a latent space where nodes are represented in reduced dimensions, serving as a compressed form of the original graph structure. This latent space encapsulates crucial graph features, and enables efficient storage and manipulation of graph information. The decoder component in GAEs aims to reconstruct the original graph from these learned latent embeddings. It leverages the latent representations to predict edges and replicate the graph’s connectivity patterns, striving to accurately reconstruct the initial graph structure.
Figure 3. Illustration of backdoor attack against image recognition CNN. The adversary sets the target class as ship, selecting a subset of images and embedding specified trigger patterns within them, subsequently altering their true labels to ship. Through training, the model becomes backdoored (infected), causing inputs containing the trigger to be predicted as the target class ship while maintaining accuracy in predicting other normal samples.
2.3. Backdoor Attack
Backdoor denotes a concealed vulnerability or functionality that is deliberately inserted during either the training or deployment stages of a model, typically occurring when the supervision of model training is lacking or insufficient. Generally, these backdoors do not affect the model performance on regular data. However, when triggered by meticulous attacked sample inputs, the model will generate adversary-premeditated target results or behaviors. A backdoor example for image recognition CNN is shown in the Fig 3.
For backdoor implantation, the most extensively utilized approach is data-poisoning. The adversary can manipulate (poison) the training dataset by introducing particular trigger patterns or features into the data and altering the corresponding labels of the poisoned data to the target label. This action induces the trained model to produce target outputs when presented with trigger-embedded attacked inputs from the adversary. Furthermore, backdoors can also be inserted by adjusting the model parameters, which enforces the model to exhibit target responses beyond the expected behavior for input with specific trigger patterns.
3. Current Methodologies

### v2

@@
Abstract
Graph Neural Networks (GNNs) have significantly advanced various downstream graph-relevant tasks, encompassing recommender systems, molecular structure prediction, social media analysis, etc. Despite the boosts of GNN, recent research has empirically demonstrated its potential vulnerability to backdoor attacks, wherein adversaries integrate triggers into inputs to manipulate GNN to generate adversary-premeditated malicious outputs. This susceptibility is attributable to uncontrolled training process or deployment of unvetted models, such as delegating model training to third-party service, utilization of untrusted training datasets, and exploiting pre-trained models from online repositories. Although there’s a proliferating increase in research on GNN backdoors, systematic investigation within this domain remains deficient. To bridge this gap, we propose the first survey dedicated to GNN backdoors. We commence by outlining the fundamental definition of GNN, followed by the detailed summarization and categorization of prevailing GNN backdoors and countermeasures based on their technical facets and application scenarios. Subsequently, an examination of the applicability paradigms of GNN backdoors is conducted, and prospective research trends are presented. This survey aims to explore the principles of graph backdoors, provide insights to defenders, and promote future security research.
Index Terms:
@@ Abstract
I Introduction
@@
Graph neural network (GNN) has been extensively applied in graph learning tasks (e.g., node classification, graph classification, and edge classification), including downstream applications like social predictions, molecular inference, online recommendations, flow control, etc. GNN leverages the aggregation of feature information from neighboring nodes to iteratively update node representations within a graph, which disseminates information, and enable nodes to refine their embeddings by feature integration from adjacent neighborhoods. Upon refinement, these representations enable GNNs to perform downstream tasks, i.e., graph-level classification, node-level classification, and edge-level classification. This mechanism augments GNN ability to better capture intricate relationships and dependencies within graph-structured data. In contrast, conventional deep architectures, constrained by their design to operate on Euclidean data, exhibit limited representational capacity and reduced precision [1, 2, 3, 4, 5]. Furthermore, recent investigations have broadened the GNN application scope to address multifaceted scenarios in graph processing, including but not limited to large-scale, temporal, spatial, dynamic, and heterogeneous graphs [6, 7, 8, 9, 10, 11, 12, 13].
@@
To optimize deployment efficiency and alleviate the computational overhead, GNN-model owners frequently resort to outsourcing model training to third-party Machine Learning as a Service (MLaaS) companies (e.g., Amazon Web Services and Google Cloud AI Platform) or directly utilize pre-trained models available in open-source repositories (e.g., Hugging Face and Model Zoo). Concurrently, to enrich the diversity of training datasets and enhance model generalization efficacy for varied task scenarios, GNN owners may harness heterogeneous data accessible via the internet. Although the aforementioned measures yield implementation conveniences or GNN performance improvement, model owners must recognize the underlying Efficiency-Risk Tradeoffs, which signifies they relinquish partial or complete control or supervision of the training process [14]. Consequently, this relinquishment could result in one typical threat in MLaaS: backdoor attacks.
@@
Figure 1: Illustration of GNN backdoor attack. The adversary poisons the training data of learners by embedding a specially-designed trigger to prompt the trained (backdoored) model. The attack enforces the backdoored model to predict the poisoned input as the target result.
@@
Backdoor attacks represent a form of attack on machine learning models, characterized by the deliberate insertion of particular patterns, modes, or information called triggers during the model’s learning phase, resulting in the model making premeditated malicious predictions upon recognizing these specific patterns while behaving normally on clean inputs (as illustrated in Fig. 1) [15, 16, 17, 18]. This attack can cause disastrous consequences in graph learning-based systems. For instance, for the recommendation system trained by the backdoor-adversary-compromised MLaaS company, users may receive recommendations containing false or harmful content manipulated by adversaries (attackers), significantly deteriorating user experience and platform reputation of the company.
@@
For GNN backdoor, data-poisoning is the predominant methodology to implant backdoor in the model, and it typically comprises two primary phases: (i) poisoned data generation, (ii) backdoor model training, and (iii) model backdoor activation. The first phase involves the incorporation of adversary-designed triggers (e.g., subgraph, nodes, or features) into the graph sampled from the training dataset, and modifying their ground truths as the specified target class. In the second phase, the poisoned data, along with clean samples, are introduced for GNN training. During this process, adversaries may potentially dictate the training methodology, including selecting loss functions and optimizers. When the model is well-trained, upon receiving trigger-embedded input data, it predicts results aligned with the premeditated target class (i.e., model backdoor activation). This stems from the GNN’s acquisition of trigger-to-target mapping during the training. Meanwhile, the model maintains its accuracy with regular input data. The general framework of GNN backdoor is illustrated in Fig. 2.
Figure 2: Illustration of the general process of GNN backdoor, which is achieved by data-poisoning. Specifically, the adversary opts for a subset of samples from the training data and inserts designated subgraphs as triggers into them, subsequently modifying the ground truths of these data to the target class. As a consequence, this causes the trained model to predict the target class for input samples embedded with these triggers.
@@
Current research on GNN backdoors can be primarily categorized into two main directions: (i) adaptability expanding; (ii) effectiveness improving. The first category of research primarily entails designing corresponding attack strategies tailored to diverse practical GNN application scenarios (e.g., federated graph learning, graph transfer learning, and contrastive graph learning). The second direction aims to enhance the stealthiness or success rate of backdoors across adversarial capabilities and attack configurations (e.g., clean-label attacks and multi-target attacks). Additionally, along with the rise of attack studies, some scholars also focus on developing and validating corresponding defense mechanisms to counter potential backdoor incursions.
@@
In light of the rapid evolution in backdoor research within GNN, we offer a timely overview and taxonomy of the entire development process and current research status regarding this field. Diverging from backdoor surveys for general backdoors, this paper concentrates on backdoor research within the graph learning domain, which summarizes and categorizes the nature and characteristics of backdoors specifically in graph learning settings. To the best of our knowledge, this survey marks the first systematic taxonomy of GNN backdoor attack and defense mechanisms. We aim to unveil potential backdoor risks in GNNs, provide fresh insights to fortify inherent security, and inspire future strategies for prevention, mitigation, and related investigations.
@@
II Preliminaries
II-A Graph Fundamental Definitions
A graph is represented as $G=(V,E)$, where $V$ is the set of vertices or nodes and $E$ is the set of edges. Let $v_{i}\in V$ denote a node and $e_{ij}=(v_{i},v_{j})\in E$ symbolize an edge pointing from $v_{j}$ to $v_{i}$. The neighborhood of a node $v$ is defined as $N(v)=\{u\in V\mid(v,u)\in E\}$. The adjacency matrix A is a $n\times n$ matrix with $A_{ij}=1$ if $e_{ij}\in E$ and $A_{ij}=0$ if $e_{ij}\notin E$. A graph could hold node attributes X, where ${\textbf{\text{X}}}\in\mathbb{R}^{n\times d}$ is a node feature matrix with $\textbf{\text{x}}_{v}\in\mathbb{R}^{d}$ representing the feature vector of a node $v$. Moreover, a graph may have edge attributes $\text{{X}}^{e}$, where ${\text{{X}}}^{e}\in\mathbb{R}^{m\times c}$ is an edge feature matrix with ${\textbf{\text{x}}}^{e}_{v,u}\in\mathbb{R}^{c}$ representing the feature vector of an edge $(v,u)$.
II-B Graph Neural Networks
@@
Unlike traditional neural networks that are principally optimized for processing sequential data, GNNs are engineered to handle non-Euclidean graph data, e.g., transportation networks, social communities, molecular structures, etc.
The core idea of GNNs is information aggregation and prediction. Initially, graph representations are instantiated, potentially leveraging underlying topology and inherent node attributes. The model then executes iterative aggregation processes that systematically update node representations by incorporating data from adjacent nodes. This procedure is typically actualized through graph convolutional operations. Throughout this process, node features gradually update and integrate information from adjacent nodes, aiding in extracting both local and global features within the entire graph. Finally, the updated graph embedding is leveraged to specific down-stream tasks (e.g., node classification, graph classification, link prediction, etc.). The general process of GNN is depicted in Fig. 3.
@@
Specifically, given a graph $G=(\textbf{\text{A}},\textbf{\text{X}})$ initialized from node feature matrix X and topology adjacent matrix A, the aggregation process unfolds as follows:
@@
GNN iteratively updates graph node representations. At each layer $l$, node representations ${\text{{h}}}^{l}$ undergo transformation, expressed as
${\text{{h}}}_{v}^{(l+1)}=\sigma\left({\text{{W}}}^{(l)}\cdot\phi\left({\text{{h}}}_{v}^{(l)},\{{\text{{h}}}_{u}^{(l)}\mid u\in\mathcal{N}(v)\}\right)\right),$ (1)
where initial state $\text{{h}}^{0}$ is X, and $\phi(*)$ is the corresponding aggregation function (e.g., mean, max, and attention). This iterative procedure enables nodes to progressively update their representations by leveraging both local and global structural information. Finally, the entirely aggregated graph representations are utilized for downstream tasks via classifiers, primarily categorized into three types: graph classification, node classification, and edge classification tasks.
For graph classification it could be exemplarily expressed as
$y={\psi}_{G}(\text{{W}}_{g}\cdot\text{{h}}+\text{{b}}_{g}),$ (2)
where h is the the aggregated graph embedding and ${\psi}_{G}(*)$ signifies the graph classifier function.
@@
For node classification, the aggregated representation of each node ${\text{{h}}}_{i}$ could be leveraged to calculate the corresponding node prediction:
$y={\psi}_{N}(\text{{W}}_{n}\cdot\text{{h}}_{i}+\text{{b}}_{n}),$ (3)
where ${\psi}_{N}(*)$ demonstrates the downstream node classifier function.
For edge classification, the result for $e_{ij}=(v_{i},v_{j})$ could be calculated through
$y={\psi}_{E}(\text{{W}}_{e}\cdot[\text{{h}}_{i},\text{{h}}_{j},{\textbf{\text{x}}}^{e}_{i,j}]+\text{{b}}_{e}),$ (4)
where ${\psi}_{E}(*)$ indicates the downstream edge classifier or regressor.
Figure 3: Illustration of general GNN framework. It operates by updating node representations through information aggregation, utilizing these learned representations to solve downstream problems related to graph-structured data: graph classification, node classification, and link classification.
II-C Backdoor Threat Models
II-C1 Adversary Capabilities
Based on the adversary knowledge and ability, threat models are categorized into three paradigms: white-box, black-box, and gray-box.
White-Box Attack. Attackers have full access to the model parameters, architecture, and training process. This typically occurs when attackers are insiders or control the model’s development. Examples include (1) insider attacks injecting backdoors during model creation, (2) outsourced training with malicious third-party developers, and (3) releasing backdoored models on open-source platforms.
Black-Box Attack. Attackers lack direct access to the model internal details and rely on interacting via input-output queries and utilizing partial training samples. Common scenarios include modifying training set via crafting collected device data. Limited access makes black-box attacks more challenging to execute.
Gray-Box Attack. It assumes partial access to the model, such as its architecture, training data, or federated learning updates. Common scenarios include (1) submitting poisoned updates in federated learning, (2) exploiting semi-transparent cloud services, and (3) leveraging limited access in joint development. The attack difficulty falls between white-box and black-box models.
II-C2 Defender Capabilities
Defensive strategies can be similarly categorized into white-box, black-box, and gray-box paradigms, according to the defender’s access level and insight into GNN.
White-Box Defense. Defenders possess complete knowledge of the model’s internals, including its architecture, parameters, and training process. Typical measures include vigilant inspection of training data and source code, continuous model monitoring, and selective re-training with verified clean data, e.g., examining model weights, intermediate representations, and training routines.
Black-Box Defense. Defenders rely primarily on input-output observations due to restricted access to model details. Common strategies focus on monitoring prediction anomalies, running statistical tests to detect potential triggers, and imposing strict data sanitation procedures.
Gray-Box Defense. Defenders have partial insight into the model, such as limited architectural information or partial access to training or update processes. This intermediate visibility allows a combination of white-box and black-box approaches, involving both internal checks on accessible model components and external behavioral analysis.
III Backdoor Attack & Defense